OpenStudy (anonymous):

You know the following about the function f. f(7) = 3 7 \leq x \leq 9 0.8 \leq f´(x) \leq 1.2 The answer is 5.4 Can someone please explain how to solve this kind of problem. @Mathematics

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

Sorry. Find the biggest value possible for f(9)

OpenStudy (anonymous):

You know the mean value theorem?

OpenStudy (anonymous):

Let me first make sure I understood your question well. We know that f(7)=3. We also know that \(0.8 \le f'(x) \le 1.2\) for \(7\le x \le9\). We're asked to find the biggest possible value for f(9).

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=1.2+%3D%3D+%28x+-+3%29%2F%289-7%29 This seems right. Thank you AnwarA

OpenStudy (anonymous):

Lol, you're welcome.